# Heavy Traffic Diffusion Limit for a Closed Queueing Network with Single-Server and Infinite-Server Stations

### v3

@@ 1.1 Literature Review
The rest of the paper is organized as follows. Section 2 introduces the model primitives and makes a sample-path construction of the queue-length processes describing the network. Section 3 articulates the heavy traffic assumption and the asymptotic regime, and states the main result of the paper (Theorem 1). Section 4 develops the key tools needed to prove the main result. Section 5 is devoted to a proof of the main result, which involves proving weak convergence of the fluid-scaled processes. Relevant notation and technical preliminaries are given shortly. Additional proofs are provided in Appendices A, B, C, and D.
@@ 2.1 Model Primitives
Jobs move through the system according to a two-level probabilistic routing structure: from single-server stations to infinite-server stations, and from infinite-server stations back to single-server stations. When a job completes service at single-server station $j\in[J]$, it is routed to infinite-server station $k\in[K]$ with probability $p_{jk}\in[0,1]$. Similarly, when a job completes service at infinite-server station $k\in[K]$, it is routed to single-server station $j\in[J]$ with probability $q_{kj}\in[0,1]$. To formalize this, we define stochastic matrices
$P\,=\,(p_{jk})\,\in\,\mathbb{R}^{J\times K},\quad Q\,=\,(q_{kj})\,\in\,\mathbb{R}^{K\times J}$ (3)
@@ 3 Heavy Traffic Assumption and Main Result
By (5)–(6), (20)–(32), and Assumption 1, it is straightforward verify that for $j\in[J]$, $k\in[K]$, and all $t\geq 0$ the following equalities hold:
@@
Also, by (28)–(29) and (33)–(34), it is straightforward to verify that for all $t\geq 0$ the following equalities hold:
@@ 5.1 Convergence of Fluid Scaled Processes
As $n\rightarrow\infty$, $(\bar{Q}^{n},I^{n},\bar{V}^{n})\Rightarrow\mathbf{0}\in D^{J+K}$
@@
Furthermore, by (14)–(15) and (28), $I^{n}$ is nondecreasing with $I^{n}(0)=0$ and satisfies
@@
Since $(\bar{\xi}^{n},\bar{\zeta}^{n})\in D^{J+K}$ pathwise, it follows from (37)–(38), (62)–(64), and Proposition 1 that $(\bar{Q}^{n},I^{n},\bar{V}^{n})=(f_{1}(\bar{\xi}^{n},\bar{\zeta}^{n}),f_{2}(\bar{\xi}^{n},\bar{\zeta}^{n}),f_{3}(\bar{\xi}^{n},\bar{\zeta}^{n}))$. Then, by Proposition 2, Lemma 2, and the continuous mapping theorem,
@@
$\,\Rightarrow\,\big(f_{1}(\mathbf{0}),f_{2}(\mathbf{0}),f_{3}(\mathbf{0})\big).$
@@ 5.2 Convergence of Diffusion Scaled Processes
Moreover, as in the proof of Lemma 3, it is straightforward to show that $\hat{I}^{n}_{j}$ is nondecreasing with $\hat{I}^{n}_{j}(0)=0$ and satisfies
[…]
@@ A.1 Existence of an element in $D_{K}^{T}$ satisfying (56) for all $t\in[0,T]$
for all $n\in\mathbb{N}$ and $k\in[K]$, where $C_{t}:=\max_{k\in[K]}\|\zeta_{k}\|_{t}+\sum_{j=1}^{J}\|\psi(\xi_{j})\|_{t}$ for $t>0$. (Note that $C_{t_{1}}\leq C_{t_{2}}$ for all $0\leq t_{1}\leq t_{2}<\infty$.) To prove this, note that when $n=1$,
@@
$\leq\,\|\zeta_{k}\|_{\delta}+\sum_{j=1}^{J}\|\psi(\xi_{j})\|_{\delta}\,\leq\,C_{\delta}.$
for all $k\in[K]$. Moreover, note that when $n\geq 2$,
@@
for all $k\in[K]$, where the first inequality follows from Whitt [25, Lemma 13.5.1] and (71). By performing a similar estimate of $\|y_{l}^{n-1}-y_{l}^{n-2}\|_{\delta}$ in the above display, it follows that
@@
for all $k\in[K]$. By continuing in this way, we obtain (73). Next, we claim that for each $k\in[K]$ that
$\|y_{k}^{n}-y_{k}^{n-1}\|_{m\delta}\,\leq\,m^{2}n^{m}(2\delta\eta JK)^{n-1}C_{m\delta}.$ (74)
for all $m,n\in\mathbb{N}$. Fixing $k\in[K]$, we consider the $n=1$ and $n\geq 2$ cases separately. When $n=1$, we have that $\|y_{k}^{1}-y_{k}^{0}\|_{m\delta}\leq C_{m\delta}\leq m^{2}C_{m\delta}$ for all $m\in\mathbb{N}$ by (73), so that (74) holds. When $n\geq 2$, we proceed by (strong) induction on $m$. The base case of $m=1$ holds immediately by (73). For the inductive step, assume that for all $n\geq 2$,
@@
for all $r\in[m]$. Thus, by (75), it follows that for $n\geq 2$,
@@
Continuing from the previous display, it follows from (73) that for $n=2$,
@@
In particular, by (74), we have that
@@ A.3 Extension to a unique element in $D^{K}$ satisfying (56) for all $t\in[0,\infty)$
$\!\!\!\!\!\!\!\!\!\!\!\!\big(y^{\infty}_{1},\dots,y^{\infty}_{K}\big)(t)\coloneqq\,\big(y^{\infty}_{1,T},\dots,y^{\infty}_{K,T}\big)(t),\quad t\in[0,T].$
for each $T>0$. We claim that $(y^{\infty}_{1},\dots,y^{\infty}_{K})$ is well-defined and is the unique element in $D^{K}$ satisfying (56) for all $t\in[0,\infty)$. To prove that it is well-defined, we must show that whenever $t\in[0,T_{1}]\cap[0,T_{2}]$, we have
@@ Appendix C Proof of Proposition 2
$\!\!\!\!\!\!\!\!\!\!\|(\xi,\zeta)\,\circ\,\lambda^{n}\,-\,(\xi^{n},\zeta^{n})\|_{T,J+K}\,\vee\,\|\dot{\lambda}^{n}\,-\,1\|_{T}\,\rightarrow\,0$ (84)
@@
see, e.g., Whitt [25, Lemma 13.5.2]. Combining these estimates, it follows from (85)–(87) that
@@
$<\,\frac{\epsilon}{e^{2\eta JKT}}\,+\,2\eta JK\!\int_{0}^{t}\!\!\max_{k\in[K]}\|y_{k}\,\circ\,\lambda^{n}\,-\,y_{k}^{n}\|_{s}\,ds$
for all $0\leq t\leq T$. By Gronwall’s inequality (see, e.g., Pang et al. [23, Lemma 4.1]) and the above displayed inequality, it follows that
$\max_{k\in[K]}\|y_{k}\,\circ\,\lambda^{n}\,-\,y^{n}_{k}\|_{t}\,<\,e^{2\eta JK(t-T)}\epsilon\,\leq\,\epsilon$ (92)

### v4

@@ 1.1 Literature Review
The rest of the paper is organized as follows. Section 2 introduces the model primitives and makes a sample-path construction of the queue-length processes describing the network. Section 3 articulates the heavy traffic assumption and the asymptotic regime, and states the main result of the paper (Theorem 1). Section 4 develops the key tools needed to prove the main result. Section 5 is devoted to a proof of the main result, which involves proving weak convergence of the fluid-scaled processes. Relevant notation and technical preliminaries are given shortly. Proofs deferred from the main text are provided in Appendices AD.
@@ 2.1 Model Primitives
Jobs move through the system according to a two-level probabilistic routing structure: from single-server stations to infinite-server stations, and from infinite-server stations back to single-server stations. When a job completes service at single-server station $j\in[J]$, it is routed to infinite-server station $k\in[K]$ with probability $p_{jk}\in[0,1]$. Similarly, when a job completes service at infinite-server station $k\in[K]$, it is routed to single-server station $j\in[J]$ with probability $q_{kj}\in[0,1]$. To formalize this, we take as given stochastic matrices
$\!\!\!\!P\,=\,(p_{jk})\,\in\,\mathbb{R}^{J\times K}\quad\text{and}\quad Q\,=\,(q_{kj})\,\in\,\mathbb{R}^{K\times J},$ (3)
@@ 3 Heavy Traffic Assumption and Main Result
By (5)–(6), (20)–(32), and Assumption 1, it is straightforward verify that for all $j\in[J]$, $k\in[K]$, and $t\geq 0$ the following equalities hold:
@@
Moreover, by (28)–(29) and (33)–(34), it is straightforward to verify that for all $j\in[J]$, $k\in[K]$, and $t\geq 0$ the following equalities hold:
@@ 5.1 Convergence of Fluid Scaled Processes
As $n\rightarrow\infty$, $(\bar{Q}^{n},I^{n},\bar{V}^{n})\Rightarrow\mathbf{0}\in D^{J+K}$.
@@
Furthermore, by (14)–(15) and (28), $I^{n}$ is nondecreasing componentwise with $I^{n}(0)=0$ and satisfies
@@
Since $(\bar{\xi}^{n},\bar{\zeta}^{n})\in D^{J+K}$ pathwise, it follows from (37)–(38), (62)–(64), and Proposition 1 that $(\bar{Q}^{n},I^{n},\bar{V}^{n})=(f_{1}(\bar{\xi}^{n},\bar{\zeta}^{n}),f_{2}(\bar{\xi}^{n},\bar{\zeta}^{n}),f_{3}(\bar{\xi}^{n},\bar{\zeta}^{n}))$. Then, by Proposition 2, Lemma 2, and the continuous mapping theorem, it follows that as $n\rightarrow\infty$,
@@
$\,\Rightarrow\,\big(f_{1}(\mathbf{0}),f_{2}(\mathbf{0}),f_{3}(\mathbf{0})\big)$
@@ 5.2 Convergence of Diffusion Scaled Processes
Moreover, as in the proof of Lemma 3, it is straightforward to show that $\hat{I}^{n}$ is nondecreasing componentwise with $\hat{I}^{n}(0)=0$ and satisfies
$\int_{0}^{\infty}\!\mathds{1}_{\{\hat{Q}_{j}^{n}(t)\,>\,0\}}\,d\hat{I}^{n}_{j}(t)\,=\,0\quad\text{for all}\quad j\in[J].$ (68)
@@ A.1 Existence of an element in $D_{K}^{T}$ satisfying (56) for all $t\in[0,T]$
for all $n\in\mathbb{N}$ and $k\in[K]$, where $C_{t}:=\max_{k\in[K]}\|\zeta_{k}\|_{t}+\sum_{j=1}^{J}\|\psi(\xi_{j})\|_{t}$ for $t>0$. (Note that $C_{t_{1}}\leq C_{t_{2}}$ for all $0\leq t_{1}\leq t_{2}<\infty$.) To prove (73), note that for $n=1$,
@@
$\leq\,\|\zeta_{k}\|_{\delta}+\sum_{j=1}^{J}\|\psi(\xi_{j})\|_{\delta}\,\leq\,C_{\delta},$
for all $k\in[K]$. Moreover, note that for $n\geq 2$,
@@
for all $k\in[K]$, where the first inequality follows from Whitt [25, Lemma 13.5.1] and (71). By performing a similar estimate for $\|y_{l}^{n-1}-y_{l}^{n-2}\|_{\delta}$ in the above display, it follows that
@@
for all $k\in[K]$. Continuing in this way, we obtain (73). Next, we claim that for each $k\in[K]$,
$\|y_{k}^{n}-y_{k}^{n-1}\|_{m\delta}\,\leq\,m^{2}n^{m}(2\delta\eta JK)^{n-1}C_{m\delta},$ (74)
for all $m,n\in\mathbb{N}$. Fixing $k\in[K]$, we consider the $n=1$ and $n\geq 2$ cases separately. When $n=1$, we have that $\|y_{k}^{1}-y_{k}^{0}\|_{m\delta}\leq C_{m\delta}\leq m^{2}C_{m\delta}$ for all $m\in\mathbb{N}$ by (73), so that (74) holds. When $n\geq 2$, we proceed by (strong) induction on $m$. The base case of $m=1$ holds immediately by (73). For the inductive step, we assume that for all $n\geq 2$,
@@
for all $r\in[m]$. Then, by (75), it follows that for $n\geq 2$,
@@
Continuing from the above display, it follows from (73) that for $n=2$,
@@
Equipped with (74), it now follows that
@@ A.3 Extension to a unique element in $D^{K}$ satisfying (56) for all $t\in[0,\infty)$
$\!\!\!\!\!\!\!\!\!\!\big(y^{\infty}_{1},\dots,y^{\infty}_{K}\big)(t)\coloneqq\,\big(y^{\infty}_{1,T},\dots,y^{\infty}_{K,T}\big)(t),\quad t\in[0,T],$
for each $T>0$. We claim that $(y^{\infty}_{1},\dots,y^{\infty}_{K})$ is well-defined and is the unique element in $D^{K}$ satisfying (56) for all $t\in[0,\infty)$. To prove that it is well-defined, we must show that whenever $t\in[0,T_{1}]\cap[0,T_{2}]$,
@@ Appendix C Proof of Proposition 2
$\!\!\!\!\!\!\!\!\!\!\|(\xi,\zeta)\,\circ\,\lambda^{n}\,-\,(\xi^{n},\zeta^{n})\|_{T,J+K}\,\vee\,\|\dot{\lambda}^{n}\,-\,1\|_{T}\,\rightarrow\,0,$ (84)
@@
see, e.g., Whitt [25, Lemma 13.5.2]. Combining these estimates, it then follows from (85)–(87) that
@@
$<\,\frac{\epsilon}{e^{2\eta JKT}}\,+\,2\eta JK\!\int_{0}^{t}\!\!\max_{k\in[K]}\|y_{k}\,\circ\,\lambda^{n}\,-\,y_{k}^{n}\|_{s}\,ds,$
for all $0\leq t\leq T$. By Gronwall’s inequality (see, e.g., Pang et al. [23, Lemma 4.1]) and the above displayed inequality, it follows that for all $n\geq N$,
$\max_{k\in[K]}\|y_{k}\,\circ\,\lambda^{n}\,-\,y^{n}_{k}\|_{t}\,<\,e^{2\eta JK(t-T)}\epsilon\,\leq\,\epsilon,$ (92)